# BLISS: A Lightweight Bilevel Influence Scoring Method for Data Selection in Language Model Pretraining

### v1

@@ 4.2 Algorithm for Updating the Proxy Model and Score Model
Now we design efficient algorithms for solving the bilevel problem (1). The lower-level problem aims to optimize the proxy model $\theta_{p}$ on the weighted training samples with the influence predicted by the score model. Note that we freeze the LLM ($\theta_{tr}$) through the process of solving the bilevel optimization problem, as the LLM is used to infer the output logits. Therefore, we perform the following update for the lower-level objective on a mini-batch of size $\mathcal{B}$:
$\theta_{p}^{t+1}$ $=\theta_{p}^{t}-\eta_{1}\nabla_{\theta_{p}}\sum_{i=1}^{\mathcal{B}}G(\theta_{p}^{t},\theta_{s}^{t};\xi_{i})$ (2)
$=\theta_{p}^{t}-\eta_{1}\sum_{i=1}^{\mathcal{B}}\Bigg(P_{i}\nabla_{\theta_{p}}\mathcal{L}(\theta_{p}^{t};\xi_{i})+\gamma\sum_{j}\nabla_{\theta_{p}}\ell_{j}(\theta_{p}^{t};\xi_{i})\log\frac{\ell_{j}(\theta_{p}^{t};\xi_{i})}{\ell_{j}(\theta_{tr}^{t};\xi_{i})}+2\lambda\theta_{p}^{t}\Bigg),$
where $\ell_{j}(\cdot)$ denotes the $j$-th logit of the output. Note that the exact computation of $P_{i}$ depends on all NN samples, which is computationally infeasible. Therefore, we approximate $P_{i}$ by replacing the full summation in the denominator with a partial summation over a smaller subset. This approximation is implemented in a distributed manner, significantly reducing the computational overhead. More details can be found in Appendix G. For the upper-level update, we take the derivative of $\Phi(\theta_{s})$ with respect to $\theta_{s}$ by chain rule, which is known as the hypergradient:
@@
$\nabla_{\theta_{s}}\widehat{\Phi}(\theta_{s}^{t})=-\sum_{i=1}^{\mathcal{B}}P_{i}\nabla_{\theta_{s}}h(\theta_{s}^{t};\xi_{i})\nabla_{\theta_{p}}\mathcal{L}(\theta_{p}^{t};\xi_{i})z^{t}+\sum_{i=1}^{\mathcal{B}}P_{i}\sum_{j=1}^{\mathcal{B}}P_{j}\nabla_{\theta_{s}}h(\theta_{s}^{t};\xi_{j})\nabla_{\theta_{p}}\mathcal{L}(\theta_{p}^{t};\xi_{i})z^{t}.$ (5)
Then the update for the upper-level variable ($\theta_{s}$) is
$\theta_{s}^{t+1}=\theta_{s}^{t}-\eta_{3}\nabla_{\theta_{s}}\widehat{\Phi}(\theta_{s}^{t}).$ (6)
@@
The detailed implementation of the algorithm is presented in Algorithm 1. The pretraining process is conducted over $R$ rounds. In each round, the algorithm performs data selection followed by LLM retraining. The training dataset is partitioned into $R$ shards. The data selection in round $r$ is conducted on $\mathcal{D}_{tr}^{r}$. The LLM resumes training from the previous round’s checkpoint and updates to $\theta_{tr}^{r}$ at the end of the $r$-th round. Similarly, the score model also continues learning throughout the process, reaching $\theta_{s}^{r}$ at the $r$-th round. It is worth noting that the proxy model ($\theta_{p}^{r}$) is reinitialized with the warm-up model at the beginning of each round. This prevents the model from overfitting to the previous round’s training data and ensures it can better capture the evolving behavior of the LLM.
@@
4:  $\theta_{p}^{0,r}=\theta_{p}^{0,0}$ # reset proxy/score parameters for the new round
5:  $\theta_{s}^{0,r}=\theta_{s}^{T,r-1}$ if $r>1$ else $\theta_{s}^{0,0}$
6:  $\theta_{tr}^{0,r}=\theta_{tr}^{Q,r-1}$ if $r>1$ else $\theta_{tr}^{0,0}$
7:  for $t=0,\dots,T-1$ do
8:   Sample $\xi_{t}^{r},\tilde{\xi}_{t}^{r},\pi_{t}^{r}\leftarrow\tilde{\mathcal{D}}_{tr}^{r}$, and sample $\zeta_{t}\leftarrow\mathcal{D}_{val}$
9:   (single/double loops) $\theta_{p}^{t+1,r}=\theta_{p}^{t,r}-\eta_{1}\nabla_{\theta_{p}}G(\theta_{p}^{t,r},\theta_{s}^{t,r};\xi_{t}^{r})$
10:   $\begin{aligned} z^{t+1,r}=\texttt{GDLS}&(\eta_{2},K,\nabla_{\theta_{p}}G(\theta_{p}^{t,r},\theta_{s}^{t,r};\tilde{\xi}_{t}^{r}),\nabla_{\theta_{p}}F(\theta_{p}^{t,r},\theta_{s}^{t,r};\zeta_{t}))\end{aligned}$
11:   $\theta_{s}^{t+1,r}=\theta_{s}^{t,r}-\eta_{3}\nabla_{\theta_{s}\theta_{p}}^{2}G(\theta_{p}^{t+1,r},\theta_{s}^{t,r};\pi_{t}^{r})z^{t+1,r}$
12:  end for
13:  Infer the influence score $\{S_{i}^{r}\mid 0\leq i\leq|\mathcal{D}_{tr}^{r}|-1\}$ on $\mathcal{D}_{tr}^{r}$ using $\theta_{s}^{T,r}$
14:  Sort $\{S_{i}^{r}\}$ in descending order and select the $20\%$ data with the highest influence scores from $\mathcal{D}_{tr}^{r}$ to form the selected data $\mathcal{D}_{s}$
15:  for $\tau=0,\dots,Q-1$ do
16:   Sample $\xi_{\tau}$ from $\mathcal{D}_{s}$.
17:   $\theta_{tr}^{\tau+1,r}=\theta_{tr}^{\tau,r}-\eta_{4}\nabla_{\theta_{tr}}\ell(\theta_{tr}^{\tau,r};\xi_{\tau})$ # pretrain the LLM
18:  end for
19:end for
@@
4:  $z_{k}=z_{k-1}$ if $k>1$ else $z_{0}$
5:  $z_{k+1}=z_{k}-\eta\big(\nabla_{\theta_{p}}^{2}g(\theta_{p})z_{k}-a\big)$
6:end for
7: Return $z_{K}$
@@ 5 Experiments
In this section, we validate the proposed bilevel influence scoring framework for pretraining data selection. Without relying on external pretrained models, we apply the bilevel optimization algorithm to train a lightweight proxy model ($\theta_{p}$) and a score model ($\theta_{s}$)) for data selection. We then pretrain a target LLM ($\theta_{tr}$), specifically Pythia-410M/1B, from scratch on a selected subset of the large-scale C4 dataset (Raffel et al., 2020), which is designed for LLM pretraining.
To assess the effectiveness of our approach, we evaluate the pretrained LLM on multiple downstream tasks and compare its performance against several baseline methods, including Random selection, DSIR (Xie et al., 2023b), SemDeDup (Abbas et al., 2023), DsDm (Engstrom et al., 2024), LESS (Xia et al., 2024), QuRating (Wettig et al., 2024), and MATES (Yu et al., 2024). We furthermore scale up our experiment to 2.8B model pretraining and achieve $1.4\%$ performance improvement over the state-of-the-art method.
@@ 5.3 Bilevel Optimization for Proxy Model and Score Model
In the Pythia-410M setting, the proxy model $\theta_{p}$ is updated with a “single-step" optimization per iteration (line 9 in Algorithm 1). However, when scaling up to larger models like Pythia-1B, we adopt a “multi-steps" update strategy for the proxy model to achieve a better lower-level solution. To demonstrate the effectiveness of bilevel optimization in training the proxy model and score model, we visulize the evolution of the lower-level training loss and upper-level validation loss during round 2 (Figure F in Appendix F) and round 5 (Figure 11(a) in Appendix F). Since the first round uses randomly selected data to warm up the LLM, our data selection algorithm is employed from the second round onward.
@@ 5.4 Evaluation Results on the Downstream Tasks
The LLM is continuously trained for 10,000 steps on the selected data in each round. After completing five rounds of training, we evaluate the zero-shot performance of Pythia-410M/1B on various downstream tasks and report the average accuracy along with the standard error for each dataset(see Table 1. Our algorithm consistently outperforms MATES and random selection methods across multiple tasks. For example on 410M setting, BLISS, compared with MATES, improves $2.4\%$ on SciQ, $0.7\%$ on ARC-E, $0.8\%$ on LogiQA, $0.2\%$ on HellaSwag, $0.4\%$ on PIQA, $0.2\%$ on WinoGrande, and $0.2\%$ on average accuracy (see Table 8). Additionally, Figure 3(a) presents the evaluation results in relation to pretraining FLOPs and training steps. BLISS consistently outperforms other baseline methods throughout the entire five-round pretraining process (with 10k steps per round). In particular, our method on 1B setting achieves a $1.7\times$ speedup in reaching the same performance as MATES, further validating the effectiveness of our data selection approach.
@@
We also explore LLaMA architecture models to validate the generalization of our method. Specifically, we use LLaMA-0.5B as the target pretraining model, and LLaMA-134M as the proxy model and score model. In each round, we first minimize the difference between the proxy model and the target model by training the proxy model toward a lower KL divergence. Then we periodically reset the proxy model to the initial state, in addition to resetting it at the beginning of each round. Table 3 presents the evaluation results compared with MATES, where BLISS exhibits strong data selection performance. At the round 33, our algorithm improves over MATES by $0.6\%$. More details are presented in Appendix B.
@@ 5.5 Computational Cost
We follows the FLOPs estimation method in Li et al. (2024) and report the total GPU FLOPs, including the pretraining, model warm-up, and data selection. Our main observation is: without relying on any external pretrained models as required in MATES, BLISS achieves higher average downstream performance while consuming fewer FLOPs. A detailed comparison of total FLOPs consumption is provided in Table 4.
With the same pretraining budget for LLM and an equivalent number of training tokens, BLISS is more efficient in data selection than MATES. The higher computational cost of MATES is due to its reliance on oracle data influence estimation, which involves computing the loss change after performing a one-step gradient descent update on an individual training sample. This process is highly time-consuming, because it requires per-sample gradient and cannot increase the batch size per GPU. In contrast, BLISS formulates data selection as a bilevel optimization problem, enabling the lightweight score model and proxy model to be trained to convergence within relatively few steps (3,000 per round). While BLISS introduces additional training steps for warming up the proxy and score models from scratch, this cost is negligible compared to the overall pretraining FLOPs.
@@
Model pretraining 6.35 78.3% Model pretraining 6.35 79.28%
@@
Data influence model inference 1.46 18.0% Data influence model inference 1.53 19.10%
@@
(a) The downstream performance of Pythia-410M/1B model w.r.t. pretraining FLOPs and steps, where the first point denotes the performance of a warm-up model trained on random data.
Table 5: Comparison of BLISS with different settings(without softmax and single level updata) over multiuple downstream datasets (410M model, 10B tokens) with 20k-step training.
@@ 6.1 Single-level versus Bilevel Optimization
In bilevel algorithm, the hyper-gradient is essential for the update of upper level parameters. To verify the effectiveness of bilevel update for the upper-level parameters, we compare bilevel update with a single update, which update $\theta_{s}$ and $\theta_{p}$ together using both training and validation data for the lower-level objective. Specifically, the upper and lower levels are reduced to a single level problem: the upper-level and lower-level parameters are updated simultaneously on validation dataset and training dataset respectively. With the same number of training steps as bilevel training, the average accuracy of single level update degrades $0.5\%$ as shown in Table 5.
@@ 6.2 KL Divergence Aligns the Proxy Model with the LLM
Our objective is to select training data that maximizes the LLM’s performance on downstream tasks. To achieve this, the proxy model must effectively represent the LLM, which we enforce by applying KL divergence loss to align their output logits. As shown in Figure 3 (Appendix C), incorporating KL divergence leads to improved performance across most downstream tasks, with a 9.3% accuracy boost on LogiQA and a 1.4% increase in average accuracy. Interestingly, while removing KL divergence results in a lower validation loss (as seen in Figure 11(b) compared to Figure F in Appendix F), it does not translate to better downstream performance. These findings highlight the importance of bridging the gap between the proxy model and the LLM to ensure effective data selection, demonstrating that a closer alignment between the two models leads to better overall performance.

### v3

@@ 4.2 Algorithm for Updating the Proxy Model and Score Model
Now we design efficient algorithms for solving the bilevel problem (1). The lower-level problem aims to optimize the proxy model $\theta_{p}$ on the weighted training samples with the influence predicted by the score model. Note that we freeze the LLM ($\theta_{tr}$) through the process of solving the bilevel optimization problem, as the LLM is used to infer the output logits. Therefore, we perform the following update for the lower-level objective on a mini-batch of size $\mathcal{B}$ (full batch is infeasible in practice):
$\theta_{p}^{t+1}$ $=\theta_{p}^{t}-\eta_{1}\nabla_{\theta_{p}}\sum_{i=1}^{\mathcal{B}}G(\theta_{p}^{t},\theta_{s}^{t};\xi_{i})$
$=\theta_{p}^{t}-\eta_{1}\sum_{i=1}^{\mathcal{B}}\Bigg(P_{i}\nabla_{\theta_{p}}\mathcal{L}(\theta_{p}^{t};\xi_{i})+\gamma\sum_{j}\nabla_{\theta_{p}}\ell_{j}(\theta_{p}^{t};\xi_{i})\log\frac{\ell_{j}(\theta_{p}^{t};\xi_{i})}{\ell_{j}(\theta_{tr};\xi_{i})}\Bigg)-2\eta_{1}\lambda\theta_{p}^{t},$ (2)
where $\ell_{j}(\cdot)$ denotes the $j$-th normalized logit of the output. Note that the exact computation of $P_{i}$ depends on all $N$ samples, which is computationally infeasible. Therefore, we approximate $P_{i}$ by replacing the full summation in the denominator with a partial summation over a smaller subset. This approximation is implemented in a distributed manner, significantly reducing the computational overhead. More details can be found in Appendix G. For the upper-level update, we take the derivative of $\Phi(\theta_{s})$ with respect to $\theta_{s}$ by chain rule, which is known as the hypergradient:
@@
$\nabla_{\theta_{s}}\widehat{\Phi}(\theta_{s}^{t})=$ $-\sum_{i=1}^{\mathcal{B}}P_{i}\nabla_{\theta_{s}}h(\theta_{s}^{t};\xi_{i})\nabla_{\theta_{p}}\mathcal{L}(\theta_{p}^{t};\xi_{i})^{T}z^{t}$ (5)
$+\sum_{i=1}^{\mathcal{B}}P_{i}\sum_{j=1}^{\mathcal{B}}P_{j}\nabla_{\theta_{s}}h(\theta_{s}^{t};\xi_{j})\nabla_{\theta_{p}}\mathcal{L}(\theta_{p}^{t};\xi_{i})^{T}z^{t}.$
@@
The detailed implementation of the algorithm is presented in Algorithm 1. In practice, we use Adam optimizer (Kingma and Ba, 2014) to update the upper-level variables, where we will update the Adam gradient with the calculated hypergradient. The pretraining process is conducted over $R$ rounds. In each round, the algorithm performs data selection followed by LLM retraining. The training dataset is partitioned into $R$ shards. The data selection in round $r$ is conducted on $\mathcal{D}_{tr}^{r}$. The LLM resumes training from the previous round’s checkpoint and updates to $\theta_{tr}^{r}$ at the end of the $r$-th round. Similarly, the score model also continues learning throughout the process, reaching $\theta_{s}^{r}$ at the $r$-th round. It is worth noting that the proxy model ($\theta_{p}^{r}$) is reinitialized with the warm-up model at the beginning of each round. This prevents the model from overfitting to the previous round’s training data and ensures it can better capture the evolving behavior of the LLM.
@@
4:  # reset proxy/score parameters for a new round
5:  $\theta_{p}^{0,r}=\theta_{p}^{0,0}$
6:  $\theta_{s}^{0,r}=\theta_{s}^{T,r-1}$ if $r\geq 1$ else $\theta_{s}^{0,0}$
7:  $\theta_{tr}^{0,r}=\theta_{tr}^{Q,r-1}$ if $r\geq 1$ else $\theta_{tr}^{0,0}$
8:  for $t=0,\dots,T-1$ do
9:   Sample $\xi_{t}^{r},\tilde{\xi}_{t}^{r},\pi_{t}^{r}\leftarrow\tilde{\mathcal{D}}_{tr}^{r}$, and sample $\zeta_{t}\leftarrow\mathcal{D}_{val}$
10:   # $L$: update the proxy model for lower-level
11:   $\theta_{p}^{t+1,r}=\theta_{p}^{t,r}-\eta_{1}\nabla_{\theta_{p}}G(\theta_{p}^{t,r},\theta_{s}^{t,r};\xi_{t}^{r})$
12:   # solve the linear system
13:   $\begin{aligned} z^{t+1,r}=\texttt{GDLS}(\eta_{2},K,\nabla_{\theta_{p}}G(\theta_{p}^{t,r},\theta_{s}^{t,r};\tilde{\xi}_{t}^{r}),\nabla_{\theta_{p}}F(\theta_{p}^{t,r},\theta_{s}^{t,r};\zeta_{t}))\end{aligned}$
14:   # UL: update the score model
15:   $\begin{aligned} \theta_{s}^{t+1,r}=\texttt{Adam}(\theta_{s}^{t,r},-\nabla_{\theta_{s}\theta_{p}}^{2}G(\theta_{p}^{t+1,r},\theta_{s}^{t,r};\pi_{t}^{r})z^{t+1,r},\eta_{3})\end{aligned}$  [^2] Adam(variable, gradient, lr) optimizer receives the current variable, its hypergradient and learning rate. Then it updates the first and second momentum, then returns the updated variable.
16:  end for
17:  Infer the influence score $\{S_{i}^{r}\mid 0\leq i\leq|\mathcal{D}_{tr}^{r}|-1\}$ on $\mathcal{D}_{tr}^{r}$ using $\theta_{s}^{T,r}$
18:  Sort $\{S_{i}^{r}\}$ in descending order and select the $20\%$ data with the highest influence scores from $\mathcal{D}_{tr}^{r}$ to form the selected data $\mathcal{D}_{s}$
19:  for $\tau=0,\dots,Q-1$ do
20:   Sample $\xi_{\tau}$ from $\mathcal{D}_{s}$.
21:   $\theta_{tr}^{\tau+1,r}=\theta_{tr}^{\tau,r}-\eta_{4}\nabla_{\theta_{tr}}\mathcal{L}(\theta_{tr}^{\tau,r};\xi_{\tau})$ # pretrain the LLM
22:  end for
23:end for
@@
4:  $z_{k+1}=z_{k}-\eta\big(\nabla_{\theta_{p}}^{2}g(\theta_{p})z_{k}-a\big)$
5:end for
6: Return $z_{K}$
@@ 5 Experiments
In this section, we validate the proposed bilevel influence scoring framework for pretraining data selection. We apply the bilevel optimization algorithm to train a lightweight proxy model ($\theta_{p}$) and a score model ($\theta_{s}$)) for data selection. We then pretrain a target LLM ($\theta_{tr}$), specifically Pythia-410M/1B, from scratch on a selected subset of the large-scale C4 dataset (Raffel et al., 2020), which is designed for LLM pretraining. we then evaluate the pretrained LLM on multiple downstream tasks and compare its performance against several baseline methods, including Random selection, DSIR (Xie et al., 2023b), SemDeDup (Abbas et al., 2023), DsDm (Engstrom et al., 2024), LESS (Xia et al., 2024), QuRating (Wettig et al., 2024), and MATES (Yu et al., 2024). We furthermore scale up our experiment to 2.8B model pretraining and achieve $1.4\%$ performance improvement over the state-of-the-art method. We also verify the the fidelity of proxy models to full-scale LLMs by domain reweighting experiment, deferred to Appendix˜J.
@@ 5.3 Bilevel Optimization for Proxy Model and Score Model
In the Pythia-410M setting, the proxy model $\theta_{p}$ is updated with a “single-step" optimization per iteration (line 9 in Algorithm 1). However, when scaling up to larger models like Pythia-1B, we adopt a “multi-steps" update strategy for the proxy model to achieve a better lower-level solution. To demonstrate the effectiveness of bilevel optimization in training the proxy model and score model, we visulize the evolution of the training loss at during round 2 (Figure 6(a) in Appendix F) and round 5 (Figure 6(b) in Appendix F). Since the first round uses randomly selected data to warm up the LLM, our data selection algorithm is employed from the second round onward.
@@ 5.4 Evaluation Results on the Downstream Tasks
The LLM is continuously trained for 10,000 steps on the selected data in each round. After completing five rounds of training, we evaluate the zero-shot performance of Pythia-410M/1B on various downstream tasks and report the average accuracy along with the standard error for each dataset(see Table 1. Our algorithm consistently outperforms MATES and random selection methods across multiple tasks. For example on 410M setting, BLISS, compared with MATES, improves $2.4\%$ on SciQ, $0.7\%$ on ARC-E, $0.8\%$ on LogiQA, $0.2\%$ on HellaSwag, $0.4\%$ on PIQA, $0.2\%$ on WinoGrande, and $0.2\%$ on average accuracy (see Table 8). Additionally, Figure 2 presents the evaluation results in relation to pretraining FLOPs and training steps. BLISS consistently outperforms other baseline methods throughout the entire five-round pretraining process (with 10k steps per round). In particular, our method on 1B setting achieves a $1.7\times$ speedup in reaching the same performance as MATES, further validating the effectiveness of our data selection approach.
@@
We also explore LLaMA architecture models to validate the generalization of our method. Specifically, we use LLaMA-0.5B as the target pretraining model, and LLaMA-134M as the proxy model and score model. In each round, we first minimize the difference between the proxy model and the target model by training the proxy model toward a lower KL divergence. Then we periodically reset the proxy model to the initial state, in addition to resetting it at the beginning of each round. Table 3 presents the evaluation results compared with MATES, where BLISS exhibits strong data selection performance. At the round $3$, our algorithm improves over MATES by $0.6\%$. More details are presented in Appendix B.
@@ 5.5 Computational Cost
We follows the FLOPs estimation method in Li et al. (2024) and report the total GPU FLOPs, including the pretraining, model warm-up, and data selection. Our main observation is: without relying on any external pretrained models as required in MATES, BLISS achieves higher average downstream performance while consuming fewer FLOPs. A detailed comparison of total FLOPs consumption is provided in Table 4, and running time/memory comparison is presented in Appendix˜H.
With the same pretraining budget for LLM and an equivalent number of training tokens, BLISS is more efficient in data selection than MATES. The higher computational cost of MATES is due to its reliance on oracle data influence estimation, which involves computing the loss change after performing a one-step gradient descent update on an individual training sample. This process is highly time-consuming, because it requires per-sample gradient and cannot increase the batch size per GPU. In contrast, BLISS formulates data selection as a bilevel optimization problem, enabling the lightweight score model and proxy model to be trained to convergence within relatively few steps, i.e., 3,000 per round (ablation study for bilevel steps is presented in Appendix˜I). While BLISS introduces additional training steps for warming up the proxy and score models from scratch, this cost is negligible compared to the overall pretraining FLOPs. One limitation of BLISS we have to acknowledge is that HVP calculation in bilevel optimization incurs higher peak memory usage, see Table˜10 for details.
@@
Model pretraining 6.35 78.3% Model pretraining 6.35 78.59%
@@
Data influence model inference 1.46 18.0% Data influence model inference 1.53 18.94%
@@
(a) Pretraining the 410M model.
(b) Pretraining the 1B model.
Figure 2: Downstream performance of Pythia-410M and Pythia-1B with respect to pretraining FLOPs and steps. The first point corresponds to a warm-up model trained on random data.
@@ 6.1 Single-level versus Bilevel Optimization
In bilevel algorithm, the hyper-gradient is essential for the update of upper level parameters. To verify the effectiveness of bilevel update for the upper-level parameters, we compare bilevel update with a single update, which update $\theta_{s}$ and $\theta_{p}$ together using both training and validation data for the lower-level objective. Specifically, the upper and lower levels are reduced to a single level problem: the upper-level and lower-level parameters are updated simultaneously on validation dataset and training dataset respectively. With the same number of training steps as bilevel training, the average accuracy of single level update degrades $0.5\%$ as shown in Table 5 in Appendix˜C.
@@ 6.2 KL Divergence Aligns the Proxy Model with the LLM
Our objective is to select training data that maximizes the LLM’s performance on downstream tasks. To achieve this, the proxy model must effectively represent the LLM, which we enforce by applying KL divergence loss to align their output logits. As shown in Figure 3 (Appendix C), incorporating KL divergence leads to improved performance across most downstream tasks, with a 9.3% accuracy boost on LogiQA and a 1.4% increase in average accuracy. Interestingly, while removing KL divergence results in a lower training loss (as seen in Figure 7 compared to Figure 6(a) in Appendix F), it does not translate to better downstream performance. These findings highlight the importance of bridging the gap between the proxy model and the LLM to ensure effective data selection, demonstrating that a closer alignment between the two models leads to better overall performance.